OpenStudy (anonymous):

find all solutions to the equation. 7 sin^2 x - 14 sin x + 2 =-5

OpenStudy (anonymous):

just apply the quadratic formula with ax^2 + bx + c = 0....

OpenStudy (anonymous):

yep, so just take the -5 to the other side, and let sin(x)=s you get 7s^2-14s+7=0 you can divide through by 7 to simplify it then just apply the quadratic formula as above just make sure in your solutions you substitute back in s=sin(x) so you'll get like s= blah which then you need to solve as it would be sin(x)=blah so x=sin^(-1) (blah)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

@sarahusher can you be more specific on the blah ?

OpenStudy (anonymous):

okay so you'll get (s^2)-2s+1=0 factorise and solve for s (s-1)(s-1)=0 s=1 (repeated) sin(x)=1 solve for x